OpenStudy (anonymous):

Please help! Find a quadratic equation with roots -1+4i and -1-4i. Show your work.

OpenStudy (jdoe0001):

do you know how to factor a quadratic equation? say \(\bf x^2+2x+1\quad ?\)

OpenStudy (jdoe0001):

well \(\bf y=x^2+2x+1\quad ?\)

OpenStudy (anonymous):

would that be (x+1)(x+1)?

OpenStudy (jdoe0001):

yes :) so what that means is that the roots of \(\bf x^2+2x+1 \implies (x+1)(x+1)\) so that means that \(\bf ((x+1)(x+1) \implies x^2+2x+1\) so you see, you'd get the polynomial BACK if you simply multiply the roots

OpenStudy (anonymous):

so multiply (-1+4i)(-1-4i)?

OpenStudy (jdoe0001):

yes \(\bf x = -1+4i \implies x+1-4i = 0 \implies (x+1-4i)=0\\ \quad \\ x = -1-4i\implies x+1+4i=0 \implies (x+1+4i)=0\\ \quad \\ \textit{so the roots are}\quad (x+1-4i)\quad and\quad (x+1+4i)\qquad so\\ \quad \\ (x+1-4i) (x+1+4i)\\ \quad \\ \textit{recall that }\qquad \color{red}{(a-b)(a+b)} = a^2-b^2\qquad thus\\ \quad \\ (x+1-4i) (x+1+4i)\implies [\color{red}{(x+1)-4i}] [\color{red}{(x+1)+(4i)}]\\ \quad \\ (x+1)^2-(4i)^2 \)

OpenStudy (jdoe0001):

and then you can expand that to get the polynomial, don't forget that \(\bf i^2 = \sqrt{(-1)^2}\implies -1\qquad thus\\ \quad \\ (4i)^2\implies 4^2i^2\implies 16(-1) \)

OpenStudy (anonymous):

okay that makes sense now thank you!

OpenStudy (jdoe0001):

yw